OpenStudy (anonymous):

Solve the given equation for x. log3 (2x-1) =2

OpenStudy (anonymous):

\[\log _{3} (2x - 1) = 2\]

OpenStudy (anonymous):

can you write this out in its equivalent exponential form?

OpenStudy (anonymous):

I am not sure what to do

OpenStudy (anonymous):

\[\large log_b M =a \] means \[\large b^a = M \]

OpenStudy (anonymous):

can you write your log equation in that second form?

OpenStudy (anonymous):

\[3^{2}=(2x-1)\] Then just solve for x?

OpenStudy (anonymous):

Thank you!